OpenStudy (haleyelizabeth2017):

Suppose you make 90% of your free throws and you attempt 3 free throws. Use the binomial theorem to calculate each probability. a) You do not make any of them. b) You only make 1 of them. c) You only make 2 of them. d) You make all of them I have the answers and all of my work, I just need it checked :)

OpenStudy (haleyelizabeth2017):

\[_{3}C_{0}(0.90^0)(0.10^3)\] \[_{3}C_{0}=\frac{3!}{0!(3-0)!=1}\]

OpenStudy (haleyelizabeth2017):

a) 0.001 Doing the same thing for the next three, I get: b) 0.027 c) 0.243 d) 0.729

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

none of them \[.9^3\]

OpenStudy (misty1212):

oops that was wrong !!

OpenStudy (misty1212):

none of them \[.1^3\]

OpenStudy (misty1212):

one of them \[3\times .1^2\times .9\]

OpenStudy (misty1212):

two of them \[3\times .1\times .9^2\]

OpenStudy (misty1212):

three of them \[.9^3\]

OpenStudy (haleyelizabeth2017):

eh?

OpenStudy (haleyelizabeth2017):

oh! okay lol

OpenStudy (haleyelizabeth2017):

Thank you!